1.Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 17,000 carburetors from the outside supplier?

2.Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $170,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 17,000 carburetors from the outside supplier?

2 Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $35 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally: 10 points 17,000 Units Per Direct materials Direct labor Variable nanufacturing overhead Fixed manufacturing overhead, traceable Fixed manufacturing overhead, all ocat ed Total cost Unit Per Year S17 $ 289,000 136,000 68,000 6 102,000 53,000 S 44 $ 748,000 eBook Hint One-third supervisory salaries, two-thirds depreciation of special equipment (no resale value). Required 1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 17,000 carburetors from the outside supplier? 2. Should the outside suppliers offer be accepted? 3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $170,000 per year. Given this new assumption, what would be financial advantage (disadvantage) of buying 17,000 carburetors from the outside supplier? 4. Given the new assumption in requirement 3, should the outside suppliers offer be accepted? Print References Complete this question by entering your answers in the tabs below Required Required 2Required 3 Required 4 Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $170,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 17,000 carburetors from the outside supplier? inancial advantage 25 Required 2 Required 4
Answer #1
1
Per unit Total
Make Buy Make Buy
Direct materials 17 289000
Direct labor 8 136000
Variable manufacturing overhead 4 68000
Fixed manufacturing overhead traceable 2 34000
Purchase cost 35 595000
Total 527000 595000
Difference in favor of making = 527000-595000 = $68000
Financial (disadvantage) $(68000)
2
Reject the offer
3
Make Buy
Total cost 527000 595000
Opportunity cost 170000
Total relevant cost 697000 595000
Difference in favor of buying = 697000-595000 = $102000
Financial advantage $102000
4
Accept the offer